OpenStudy (anonymous):

factor 25x^2y^2+95xy^2-150y^2

ganeshie8 (ganeshie8):

start by pulling out the GCF

ganeshie8 (ganeshie8):

you have 3 terms there right ? wats the common thing in all 3 terms ?

OpenStudy (anonymous):

y^2

ganeshie8 (ganeshie8):

Yup ! factor that out first

ganeshie8 (ganeshie8):

\( 25x^2y^2+95xy^2-150y^2 \) \(y^2( 25x^2+95x-150) \)

ganeshie8 (ganeshie8):

also looks like we can take out 5 also

OpenStudy (anonymous):

y(y+?)(y-?) something like that right?

ganeshie8 (ganeshie8):

\( 25x^2y^2+95xy^2-150y^2 \) \(y^2( 25x^2+95x-150) \) \(5y^2( 5x^2+19x-30) \)

OpenStudy (anonymous):

how did u get that?

ganeshie8 (ganeshie8):

25 = 5*5 95 = 5*19 150 = 5*30

ganeshie8 (ganeshie8):

ive just factored out the 5, like we factored the y^2 before

OpenStudy (anonymous):

it was wrong ;(

ganeshie8 (ganeshie8):

we're not done yet :o

OpenStudy (anonymous):

ooooh lol

ganeshie8 (ganeshie8):

:) \( 25x^2y^2+95xy^2-150y^2 \) \(y^2( 25x^2+95x-150) \) \(5y^2( 5x^2+19x-30) \)

ganeshie8 (ganeshie8):

fine so far right ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

i really appreciate ur help btw. :)

ganeshie8 (ganeshie8):

now, think of two numbers such that : sum = 19 product = 5(-30) = -150

ganeshie8 (ganeshie8):

25 - 6 = 19 25(-6) = -150 so these are the magic numbers here

ganeshie8 (ganeshie8):

\( 25x^2y^2+95xy^2-150y^2 \) \(y^2( 25x^2+95x-150) \) \(5y^2( 5x^2+19x-30) \) \(5y^2( 5x^2+25x - 6x -30) \) \(5y^2( 5x(x+5) - 6(x+5)) \) \(5y^2( (x+5)(5x-6)) \) \(5y^2 (x+5)(5x-6) \)

ganeshie8 (ganeshie8):

we're done. looks that was mouthful :o see if it makes some sense :)

OpenStudy (anonymous):

it does ! thank u soooo much!

ganeshie8 (ganeshie8):

glad to hear ! np you're wlc :)